(period of orbit)^{2} = K (radius of orbit)^{3}

\frac{(period of Mars)^{2}}{(period of Earth)^{2}} = \frac{K (radius of Mars)^{3}}{K (radius of Earth)^{3}}

\frac{(686)^{2}}{(365)^{2}}  = \frac{(radius of Mars)^{3}}{(radius of Earth)^{3}}

3.532  = \frac{(radius of Mars)^{3}}{(radius of Earth)^{3}}